CONVERT the Following Length: 75.225x10¹³ nm TO mm


1
Expert's answer
2021-09-09T10:51:51-0400

1 nanometer is 1/1,000,000,000 of 1 meter. 1 meter is 1000 mm, which means that 1 nanometer is 1/1,000,000,000*1000=1/1,000,000 mm. So:


"L=75.255·10^{13}·\frac{1}{1,000,000}=75.255·10^7\text{ mm}."